two thermos bottles were sold on rs.1920 together bearing 20% profit in the first and 20% loss next. find the selling price of each thermos bottle

Grade:8

1 Answers

Abhishek Srivastava
2 years ago
Let the cost price of each thermos bottle be x.
One bottle is sold at 20% profit, therefore Selling price = C.P. + C.P. x 20/100 = x + x/5 = 6x/5
Another bottle is sold at 20%loss, therefore Selling price = C.P. - C.P. x 20/100 = x - x/5 = 4x/5
Total selling price = 6x/5 + 4x/5 = 2x
Given, total selling price = Rs. 1920
Therefore, 2x = Rs. 1920
=> x = Rs. 960
Hence, 
Selling price of first bottle = 6x/5 = Rs. 1152
Selling price of second bottle = 4x/5 = Rs. 768